# Query Expansion in the Age of Pre-trained and Large Language Models: A Comprehensive Survey

### v1

@@ Abstract.
Modern information retrieval (IR) must bridge short, ambiguous queries and ever more diverse, rapidly evolving corpora. Query Expansion (QE) remains a key mechanism for mitigating vocabulary mismatch, but the design space has shifted markedly with pre-trained language models (PLMs) and large language models (LLMs). This survey synthesizes the field from three angles: (i) a four-dimensional framework of query expansion - from the point of injection (explicit vs. implicit QE), through grounding and interaction (knowledge bases, model-internal capabilities, multi-turn retrieval) and learning alignment, to knowledge graph-based argumentation; (ii) a model-centric taxonomy spanning encoder-only, encoder-decoder, decoder-only, instruction-tuned, and domain/multilingual variants, highlighting their characteristic affordances for QE (contextual disambiguation, controllable generation, zero-/few-shot reasoning); and (iii) practice-oriented guidance on where and how neural QE helps in first-stage retrieval, multi-query fusion, re-ranking, and retrieval-augmented generation (RAG). We compare traditional query expansion with PLM/LLM-based methods across seven key aspects, and we map applications across web search, biomedicine, e-commerce, open-domain QA/RAG, conversational and code search, and cross-lingual settings. The review distills design grounding and interaction, alignment/distillation (SFT/PEFT/DPO), and KG constraints - as robust remedies to topic drift and hallucination. We conclude with an agenda on quality control, cost-aware invocation, domain/temporal adaptation, evaluation beyond end-task metrics, and fairness/privacy. Collectively, these insights provide a principled blueprint for selecting and combining QE techniques under real-world constraints.
@@ 1. Introduction
Sec. 2 revisits foundations and traditional QE methods. Sec. 3 overviews pre-trained model families. Sec. 4 organizes Query Expansion techniques into four key dimensions: Point of Injection(implicit/embedding, selection-based explicit), Grounding and Interaction(integration of external knowledge and feedback), Learning and Alignment(training methods like SFT, PEFT, DPO), and KG-Augmented. Sec. 5 provides a comparative analysis of traditional vs. neural QE. Sec. 6 surveys application domains and use cases. Sec. 7 discusses open challenges and future directions.
{forest} forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=draw-leaf, rounded corners, align=left, text centered, minimum width=5em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=1.2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=0text width=10.5em,fill=fill_0,draw=draw_0,font=,, where level=1text width=10.5em,fill=fill_1,draw=draw_1,font=,, where level=2text width=14em,fill=fill_2,draw=draw_2,font=,, where level=3font=,, [ Query
Expansion, minimum height=2.8em[ Traditional
Query Expansion [ Global Analysis: Corpus-Wide
and External Knowledge [ Corpus-Wide Statistical Methods [ Term Co-occurrence and Clustering:  Minker et al. (1972),  Willett (1988),
 Peat and Willett (1991), , leaf, text width=27.8em ] [ Global Similarity Thesaurus:  Qiu and Frei (1993) , leaf, text width=27.8em ] ] [ Pre-defined Knowledge Resource [ General-purpose Lexicons:  Voorhees (1993) , leaf, text width=27.8em ] [ Domain Ontologies:  Bhogal et al. (2007),  Lu et al. (2009),  Aronson and Rindflesch (1997),
 Fu et al. (2005) , leaf, text width=27.8em ] ] ] [ Local Analysis:
Query-Dependent Evidence [ PRF/RF: Rocchio Jr (1971), Robertson and Jones (1976) , leaf, text width=27.8em ] [ LCA: Xu and Croft (2000, 2017) , leaf, text width=27.8em ] ] [ Query Logs and User Behavior [Mining reformulations or Click-based semantic proximity:  Wen et al. (2002),
 Cui et al. (2002),  Cui et al. (2003),  Baeza-Yates et al. (2004),  Yin et al. (2009) , leaf, text width=27.8em ] ] ] [ Query Expansion of
@@
PLMs/LLMs [ Pre-trained Model Landscape [ Encoder-only: BERT (Devlin et al., 2019), RoBERT (Liu et al., 2019b) , leaf, text width=27.8em ] [ Encoder–Decoder: BART (Lewis et al., 2019), T5 (Raffel et al., 2019) , leaf, text width=27.8em ] [ Decoder-only LLMs GPT-3/4 (Floridi and Chiriatti, 2020; Achiam et al., 2023), PaLM (Anil et al., 2023), LLaMA-family , leaf, text width=27.8em ] [ Instruction-tuned Models: InstructGPT, FLAN-T5, LLaMA-2 Chat,(Ouyang et al., 2022; Wei et al., 2021; Touvron et al., 2023) , leaf, text width=27.8em ] [ Domain-specific / Multilingual: BioBERT(Lee et al., 2019), UMLS-BERT(Michalopoulos et al., 2020), SciBERT(Beltagy et al., 2019),
multilingual BERT/RoBERTa families , leaf, text width=27.8em ] ] [ PLM/LLM-Driven QE Techniques [ Point of injection [ Implicit Embedding-based QE: ANCE-PRF (Yu et al., 2021), ColBERT-PRF (Wang et al., 2023b), Eclipse (D’Erasmo et al., 2025),
@@
QB-PRF (Zhang et al., 2024a), LLM-VPRF (Li et al., 2025c) ,leaf, text width=27.8em ] [ Selection-based Explicit QE: CEQE (Naseri et al., 2021), SQET (Naseri et al., 2022), BERT-QE (Zheng et al., 2020), CQED (Khader and Ensan, 2023),
PQEWC (Bassani et al., 2023) ,leaf, text width=27.8em ] ] [ Grounding and interaction [ Zero-Grounding, Non-Interactive QE: LLM-Driven Single-Stage Expansion:
Query2Doc (Wang et al., 2023c), CoT-QE (Jagerman et al., 2023), GAR (Mao et al., 2020), GRF (Mackie et al., 2023b), HyDE (Gao et al., 2023), Exp4Fuse (Liu and Zhang, 2025),
Contextual clue sampling and fusion (Liu et al., 2022), SEAL (Bevilacqua et al., 2022) ,leaf, text width=31em ] [ Grounding-Only, Non-Interactive QE: Corpus-Evidence Anchored Single-Pass Expansion:
MILL (Jia et al., 2023), AGR (Chen et al., 2024), EAR (Chuang et al., 2023), GenPRF (Wang et al., 2023a), CSQE (Lei et al., 2024), MUGI (Zhang et al., 2024b), PromptPRF (Li et al., 2025b),
FGQE (Jaenich et al., 2025) ,leaf, text width=31em ] [ Grounding-Aware Interactive QE: Multi-Round Retrieve-Expand Loops:
InteR (Feng et al., 2023), ProQE (Rashid et al., 2024), LameR (Shen et al., 2023), ThinkQE (Lei et al., 2025) ,leaf, text width=31em ] ] [ Learning and Alignment for QE with PLMs/LLMs [ SoftQE (Pimpalkhute et al., 2024), RADCoT (Lee et al., 2024), ExpandR (Yao et al., 2025), AQE (Yang et al., 2025) ,leaf, text width=18em ] ] [ KG-augmented Query Expansion [ Entity-based Expansion via Knowledge Graph Retrieval: KGQE (Perna, 2025),
CL-KGQE (Rahman et al., 2019) ,leaf, text width=25em ] [ Entity-based Expansion via Hybrid KG and Document Graph: KAR (Xia et al., 2025), QSKG (Mackie and Dalton, 2022) ,leaf, text width=29.3em ] ] ] [ Application Domains[ Web Search Engines: EQE (Zhang et al., 2023) , leaf, text width=27.8em ] [ Biomedical Information Retrieval: Kelly (2021), Khader et al. (2022),
Ateia and Kruschwitz (2025), Niu et al. (2024) , leaf, text width=27.8em ] [ E-commerce Search: BEQUE (Peng et al., 2024),  Wang et al. (2024b),  Zhang et al. (2022) ,leaf, text width=27.8em ] [ Cross-Lingual and Multilingual Search: Rajaei et al. (2024) , leaf, text width=27.8em ] [ Open-Domain Question Answering: GAR (Mao et al., 2020), EAR (Chuang et al., 2023), AQE (Yang et al., 2025), MILL (Jia et al., 2023),
AGR (Chen et al., 2024), HTGQE (Zhu et al., 2023), TDPR (Li et al., 2025a) , leaf, text width=27.8em ] [ Retrieval-Augmented Generation (RAG): QE-RAGpro (Hu et al., [n. d.]), QOQA (Koo et al., 2024),
KG-Infused RAG (Wu et al., 2025),  Rajaei et al. (2024) , leaf, text width=27.8em ] [ Conversational Search: PRL (Mo et al., 2023), MISE (Kumar and Callan, 2020), QuReTeC (Voskarides et al., 2020) , leaf, text width=27.8em ] [ Code Search: GACR (Li et al., 2022), SSQR (Mao et al., 2023), SG-BERT/GPT2 (Liu, 2023), ECSQE (Bibi et al., 2025) , leaf, text width=27.8em ] ] ] ]
@@ 7. Future Directions and Open Challenges
Although PLM/LLM-driven QE has delivered strong gains across web, biomedical, e-commerce, ODQA, RAG, conversational and code search settings, our synthesis also reveals recurring failure modes and systems bottlenecks. We group forward-looking directions by the concrete issues they address, and anchor each to empirical evidence summarized in Sections 26.
7.1. Reliability on unfamiliar and ambiguous queries
Empirical basis. When domain knowledge is sparse or the query is highly polysemous, LLM expansions may hallucinate entities or collapse onto a single popular sense, reducing facet coverage and recall (Abe et al., 2025). HyDE/Q2D-style expansions can be fluent yet off-topic in low-overlap domains (see Section 4 and the RAG discussion).
Opportunities. (i) Retrieval-augmented prompting that injects corpus-specific evidence pre- or in-generation (Mackie et al., 2023b; Lei et al., 2024). (ii) Facet-aware diversification via multi-intent expansion, multi-query issuing and rank fusion, with ambiguity detectors that trigger sense-splitting pipelines; complemented by expansion-level quality metrics (e.g., Expansion Gain@k, Factuality Score) to detect and quantify hallucinations in unfamiliar or polysemous query settings. (iii) Difficulty-aware routing that classifies queries as ”familiar” (use prompt-only QE) or ”unfamiliar” (use feedback-grounded QE) via query-quality predictors (Chen et al., 2025).
7.2. Knowledge leakage and generalization
Empirical basis. On zero-shot verification and ODQA, gains sometimes coincide with expansions that paraphrase gold passages, indicating leakage from LLM pretraining (Yoon et al., 2025).
Opportunities. (i) Leakage-controlled benchmarks with temporal splits and held-out entities, plus reporting of evidence originality. (ii) Detectors that measure n-gram or embedding overlap against evaluation corpora and pretraining proxies. (iii) Dynamic grounding that biases generation toward fresh external sources or KGs rather than static parametric memory.
7.3. Controlling expansion quality (safety, drift, usefulness)
Empirical basis. Even strong QE can degrade precision through subtle drift; automated filtering remains challenging (Weller et al., 2023).
Opportunities. (i) Pre-use vetting with lightweight judges for relevance, faithfulness and verifiability, rejecting expansions before retrieval. (ii) Counterfactual checks such as drop-term and swap-sense ablations to ensure monotonic utility of each expansion unit. (iii) Human-in-the-loop editing in high-stakes domains, in the spirit of editable DSL queries in biomedical search.
7.4. Efficiency, scalability, and selective invocation
Empirical basis. Prompt-only QE (e.g., Q2D) often adds seconds-level latency per query; dense+LLM stacks are costly. Distillation shows large savings with limited loss (Pimpalkhute et al., 2024; Lee et al., 2024).
Opportunities. (i) Distillation and PEFT: compress LLM-QE into student encoders (representation-level SoftQE; reasoning-level RADCoT) for online budgets. (ii) Selective invocation: classifiers route only ambiguous/short/tail queries to expensive QE (Chen et al., 2025). (iii) Caching and reuse: query-semantic clustering with shared expansions, TTLs by domain, and offline precomputation for tail traffic.
@@ 7.5. Dynamic domains and continual adaptation
Empirical basis. Static PLMs/LLMs lag emerging terms and events; classical PRF is corpus-specific but brittle without guards.
Opportunities. (i) Continual retrieval-augmented QE that blends LLM priors with on-the-fly corpus/KG snippets (Mackie et al., 2023b; Lei et al., 2024). (ii) Low-shot domain transfer with instruction packs and adapter-based PEFT to learn domain-specific expansion styles. (iii) Drift detectors that monitor expansion–click divergence and term novelty, auto-switching to conservative selection methods such as CEQE/SQET when risk is high (Naseri et al., 2021, 2022).
7.6. Evaluation beyond end-to-end IR metrics
Empirical basis. MAP/nDCG reflect end impact but hide why QE helped or hurt.
Opportunities. (i) Expansion-level diagnostics: Expansion Gain@$k$ (fraction of new relevant docs attributable to QE), Precision of Added Terms (share of QE-induced clicks that are relevant), and Facet Coverage (number of gold facets touched by multi-query QE). (ii) User-centric measures: satisfaction on ambiguous tasks and edit-acceptance rates in interactive QE. (iii) RAG-specific groundedness/attribution under QE-augmented retrieval.
7.7. Fairness, robustness, and governance
Empirical basis. Exposure can skew toward head facets/providers; fairness-aware QE is promising but early-stage (Jaenich et al., 2025).
Opportunities. (i) Fairness-aware generation optimizing exposure parity under minimal nDCG loss (e.g., AWRF-style). (ii) Robustness audits across tail intents, languages and domain shifts. (iii) Privacy and transparency: disclose expansion edits; prefer on-device or anonymized QE when logs are sensitive.
7.8. Integration patterns with neural IR and RAG
Empirical basis. QE helps at different pipeline points: first-stage recall (GAR/EAR/AGR), multi-query fusion, re-ranking (BERT-QE), and RAG query optimization (QOQA, KG-infused RAG) (Mao et al., 2020; Chuang et al., 2023; Chen et al., 2024; Koo et al., 2024; Wu et al., 2025).
Opportunities. (i) Design guides that choose explicit (term/text), implicit (embedding) or hybrid QE by corpus lexicality and compute budget. (ii) Constrained generation using corpus-valid substrings as actionable keys for latency-critical search (Bevilacqua et al., 2022). (iii) Multimodal and cross-lingual QE leveraging multilingual/domain encoders and back-translation fusion.
7.9. From research artifacts to reproducible systems
Empirical basis. Results are sensitive to prompts, decoding, PRF pools, and retrieval backends; leakage and cost reporting are often omitted.
Checklist. Release prompts/in-context examples/decoding parameters; report compute, latency and cost per 1k queries for online/offline modes; include leakage controls and temporal splits; ablate selection vs. generation vs. hybrid; ship reproducible GRF/PRF/QE fusion scripts.
Takeaway
Near term, the frontier is reliable, economical, and controllable QE: diversification-aware generation grounded in corpus/KG evidence, selective invocation with strong filters, and distilled/PEFT students for production. Longer term, QE will co-evolve with neural IR and RAG through constrained, leakage-aware expansion for recall, interpretable and fairness-aware exposure, and diagnostics that make expansion effects explicit.
@@ 8. Conclusion
This survey traced query expansion from classical corpus- and lexicon-driven methods to contemporary PLM/LLM paradigms. We organized the space along four analytical dimensions: point of injection, which distinguishes explicit versus implicit expansion within retrieval or generation pipelines; grounding and interaction, which captures query anchoring to knowledge bases, multi-turn retrieval, and model-internal capabilities; learning alignment, which addresses how models acquire and optimize expansion strategies; and knowledge graph–based argumentation, which leverages structured knowledge and multi-hop reasoning to support interpretable expansion, pipeline position, and application arenas. Across domains, empirical findings show that effective query expansion extends beyond pre-retrieval reformulation to multiple stages of IR pipelines, including first-stage retrieval, re-ranking, and retrieval-augmented generation. The four dimensions provide complementary levers: the injection point guides impact, grounding, and interaction ensure relevance, learning alignment supports domain/task adaptation, and knowledge graph–based argumentation enables interpretable, structured reasoning. Together, they offer a principled framework for designing and combining QE techniques under practical constraints, while pointing to future challenges in quality control, cost-aware deployment, adaptation, and fairness-aware evaluation.
We also distilled practical guidance: choose explicit vs. implicit vs. hybrid QE by corpus lexicality and latency budget; prefer feedback-grounded or constrained generation in safety-critical settings; and use selective invocation plus distillation to control cost. Finally, we outlined open challenges with concrete, evidence-backed directions, including reliability under ambiguity, leakage-aware evaluation, efficiency at scale, continual adaptation, fairness and governance, and reproducible system building.
Overall, the path forward is a synthesis of old and new: combine human-curated knowledge (ontologies, user behavior) with machine intelligence (instruction-following, reasoning, alignment) to deliver robust, transparent, and efficient expansion that consistently improves retrieval and downstream generation.

### v2

@@ Abstract.
Modern information retrieval (IR) must reconcile short, ambiguous queries with increasingly diverse and dynamic corpora. Query expansion (QE) remains central to alleviating vocabulary mismatch, yet the design space has shifted with pre-trained and large language models (PLMs, LLMs). In this survey, we organize recent work along four complementary dimensions: the point of injection (implicit/embedding vs. selection-based explicit), grounding and interaction (from zero-grounding prompts to multi-round retrieve–expand loops), learning and alignment (SFT/PEFT/DPO), and knowledge-graph integration. A model-centric taxonomy is also outlined, spanning encoder-only, encoder–decoder, decoder-only, instruction-tuned, and domain or multilingual variants, with affordances for QE such as contextual disambiguation, controllable generation, and zero-shot or few-shot reasoning. Practice-oriented guidance specifies where neural QE helps most: first-stage retrieval, multi-query fusion, re-ranking, and retrieval-augmented generation (RAG). The survey compares traditional and neural QE across seven aspects and maps applications in web search, biomedicine, e-commerce, open-domain question answering/RAG, conversational and code search, and cross-lingual settings. The survey concludes with an agenda focused on reliable, safe, efficient, and adaptive QE, offering a principled blueprint for deploying and combining techniques under real-world constraints.
@@ 1. Introduction
Section 2 revisits foundations and traditional QE methods. Section 3 overviews pre-trained model families. Section 4 organizes Query Expansion techniques into four key dimensions: Point of Injection(implicit/embedding, selection-based explicit), Grounding and Interaction(integration of external knowledge and feedback), Learning and Alignment(training methods like SFT, PEFT, DPO), and KG-Augmented. Section 5 provides a comparative analysis of traditional vs. neural QE. Section 6 surveys application domains and use cases. Section 7 discusses open challenges and future directions.
[…]
Corpus-Wide Statistical Methods
Term Co-occurrence and Clustering:  Minker et al. (1972),  Willett (1988), Peat and Willett (1991), Global Similarity Thesaurus:  Qiu and Frei (1993)
Pre-defined Knowledge Resource
@@
General-purpose Lexicons:  Voorhees (1993) Domain Ontologies:  Bhogal et al. (2007),  Lu et al. (2009),  Aronson and Rindflesch (1997), Fu et al. (2005) Local Analysis: Query-Dependent Evidence PRF/RF: Rocchio Jr (1971), Robertson and Jones (1976) LCA: Xu and Croft (2000, 2017) Query Logs and User Behavior Mining reformulations or Click-based semantic proximity:  Wen et al. (2002), Cui et al. (2002),  Cui et al. (2003),  Baeza-Yates et al. (2004),  Yin et al. (2009) Query Expansion of PLMs/LLMs Pre-trained Model Landscape Encoder-only: BERT (Devlin et al., 2019), RoBERT (Liu et al., 2019b) Encoder–Decoder: BART (Lewis et al., 2019), T5 (Raffel et al., 2019) Decoder-only LLMs GPT-3/4 (Floridi and Chiriatti, 2020; Achiam et al., 2023), PaLM (Anil et al., 2023), LLaMA-family Instruction-tuned Models: InstructGPT, FLAN-T5, LLaMA-2 Chat,(Ouyang et al., 2022; Wei et al., 2021; Touvron et al., 2023) Domain-specific / Multilingual: BioBERT(Lee et al., 2019), UMLS-BERT(Michalopoulos et al., 2020), SciBERT(Beltagy et al., 2019), multilingual BERT/RoBERTa families PLM/LLM-Driven QE Techniques
Point of injection
@@
Implicit Embedding-based QE: ANCE-PRF (Yu et al., 2021), ColBERT-PRF (Wang et al., 2023b), Eclipse (D’Erasmo et al., 2025), QB-PRF (Zhang et al., 2024a), LLM-VPRF (Li et al., 2025c) Selection-based Explicit QE: CEQE (Naseri et al., 2021), SQET (Naseri et al., 2022), BERT-QE (Zheng et al., 2020), CQED (Khader and Ensan, 2023), PQEWC (Bassani et al., 2023)
Grounding and interaction
Zero-Grounding, Non-Interactive QE: LLM-Driven Single-Stage Expansion: Query2Doc (Wang et al., 2023c), CoT-QE (Jagerman et al., 2023), GAR (Mao et al., 2020), GRF (Mackie et al., 2023b), HyDE (Gao et al., 2023), Exp4Fuse (Liu and Zhang, 2025), Contextual clue sampling and fusion (Liu et al., 2022), SEAL (Bevilacqua et al., 2022) Grounding-Only, Non-Interactive QE: Corpus-Evidence Anchored Single-Pass Expansion: MILL (Jia et al., 2023), AGR (Chen et al., 2024), EAR (Chuang et al., 2023), GenPRF (Wang et al., 2023a), CSQE (Lei et al., 2024), MUGI (Zhang et al., 2024b), PromptPRF (Li et al., 2025b), FGQE (Jaenich et al., 2025) Grounding-Aware Interactive QE: Multi-Round Retrieve-Expand Loops: InteR (Feng et al., 2023), ProQE (Rashid et al., 2024), LameR (Shen et al., 2023), ThinkQE (Lei et al., 2025)
Learning and Alignment for QE with PLMs/LLMs
SoftQE (Pimpalkhute et al., 2024), RADCoT (Lee et al., 2024), ExpandR (Yao et al., 2025), AQE (Yang et al., 2025)
KG-augmented Query Expansion
Entity-based Expansion via Knowledge Graph Retrieval: KGQE (Perna, 2025), CL-KGQE (Rahman et al., 2019) Entity-based Expansion via Hybrid KG and Document Graph: KAR (Xia et al., 2025), QSKG (Mackie and Dalton, 2022) Application Domains Web Search Engines: EQE (Zhang et al., 2023) Biomedical Information Retrieval: Kelly (2021), Khader et al. (2022), Ateia and Kruschwitz (2025), Niu et al. (2024) E-commerce Search: BEQUE (Peng et al., 2024),  Wang et al. (2024b),  Zhang et al. (2022) Cross-Lingual and Multilingual Search: Rajaei et al. (2024) Open-Domain Question Answering: GAR (Mao et al., 2020), EAR (Chuang et al., 2023), AQE (Yang et al., 2025), MILL (Jia et al., 2023), AGR (Chen et al., 2024), HTGQE (Zhu et al., 2023), TDPR (Li et al., 2025a) Retrieval-Augmented Generation (RAG): QE-RAGpro (Hu et al., [n. d.]), QOQA (Koo et al., 2024), KG-Infused RAG (Wu et al., 2025),  Rajaei et al. (2024) Conversational Search: PRL (Mo et al., 2023), MISE (Kumar and Callan, 2020), QuReTeC (Voskarides et al., 2020) Code Search: GACR (Li et al., 2022), SSQR (Mao et al., 2023), SG-BERT/GPT2 (Liu, 2023), ECSQE (Bibi et al., 2025)
@@ 7. Future Directions and Open Challenges
Despite notable gains, PLM/LLM-driven QE still faces recurring weaknesses. We outline four forward-looking challenges aligned with the four dimensions of Section 4.
7.1. Reliability under unfamiliar and ambiguous queries
Evidence. In sparse or long-tail domains, LLM expansions may hallucinate entities or collapse onto a dominant sense, reducing recall and facet coverage (Abe et al., 2025). HyDE/Q2D-style pseudo-documents are often fluent but off-topic when corpus overlap is low. Directions. Retrieval-augmented prompting can inject corpus-specific evidence (Mackie et al., 2023b; Lei et al., 2024). Facet-aware diversification (multi-intent expansion, multi-query fusion) together with ambiguity detectors can trigger sense-splitting. Difficulty-aware routing may classify queries as “familiar” (prompt-only) or “unfamiliar” (feedback-grounded) (Chen et al., 2025).
7.2. Controlling expansion quality and safety
Evidence. Even strong QE can degrade precision through topic drift, while automated filtering remains unreliable (Weller et al., 2023). Directions. Lightweight judges for relevance, faithfulness, and verifiability can vet expansions pre-use. Counterfactual checks (drop-term or swap-sense) help ensure monotonic utility. In high-stakes domains, interactive editing may keep expansions faithful and auditable.
7.3. Efficiency and selective invocation
Evidence. Generative QE introduces seconds-level latency and high compute cost. Distillation shows large savings with limited loss (Pimpalkhute et al., 2024; Lee et al., 2024). Directions. Distillation and PEFT can compress LLM-QE into efficient student models. Selective invocation routes only ambiguous or tail queries to expensive QE (Chen et al., 2025), while caching and reuse exploit query-semantic clusters to amortize cost.
@@ 7.4. Dynamic domains and continual adaptation
Evidence. Static PLMs/LLMs lag behind emerging terms and events; classical PRF is corpus-specific but brittle. Directions. Continual retrieval-augmented QE blends LLM priors with on-the-fly corpus or KG evidence (Mackie et al., 2023b; Lei et al., 2024). Low-shot transfer with adapters or instruction packs supports domain-specific styles. Drift detectors can monitor expansion–click divergence and switch to conservative selection-based QE when risk is high (Naseri et al., 2021, 2022).
Takeaway.
The near-term frontier is reliable, controllable, and economical QE: corpus-grounded expansions for ambiguous queries, lightweight filters for safety, and distilled students for efficiency. Longer term, QE should adapt continually to dynamic domains while co-evolving with neural IR and RAG, guided by evaluation and governance that make expansion effects explicit.
@@ 8. Conclusion
This survey traces the evolution of query expansion from classical corpus- and lexicon-driven methods to contemporary PLM/LLM paradigms. It organized the field along four complementary dimensions: point of injection, grounding and interaction, learning and alignment, and knowledge-graph integration. These dimensions clarify how QE can be designed and combined across retrieval pipelines, from first-stage recall to re-ranking and retrieval-augmented generation.
The survey also provides a model-centric taxonomy and distilled practical guidance for deployment, including when to select explicit, implicit, or hybrid QE, how to apply feedback-grounded or constrained generation in safety-critical settings, and how to control cost through selective invocation and distillation.
Finally, the survey highlights four open challenges that define a forward agenda: improving reliability under ambiguity, strengthening quality and safety control, achieving efficiency at scale, and enabling continual adaptation to dynamic domains. Addressing these challenges will be key to advancing QE research and practice under real-world constraints.